OpenStudy (anonymous):

The Community Recreation Center is developing plans for a new sports facility. Community members can submit suggestions for the new facility, along with basic scale drawings of their ideas. Elena wants to include a new 15- by 30-meter basketball court in the athletic center. She is submitting a scale drawing on an 8.5- by 11-inch sheet of paper. Which scale should Elena use to create as large a drawing as possible on the paper?

OpenStudy (anonymous):

A. 1/2 in. = 1 m B. 5/16 in. = 1 m C. 1/4 in. = 1 m D. 3/8 in. = 1 m